OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. (5, 3) and (6, 8)

jimthompson5910 (jim_thompson5910):

Hint: Use the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Use the formula: y2-y1=m(x2-x10

OpenStudy (anonymous):

\[8-3=m(6-5)\] so u get 5=m m is the slope